OpenStudy (anonymous):

Which inequality models this problem? The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 2w + 2 • (5w) ≤ 96 B. 2w + 2 • (5w) ≥ 96 C. 2w + 2 • (5w) < 96 D. 2w + 2 • (5w) > 96

OpenStudy (anonymous):

@campbell_st @Destinymasha @tkhunny

OpenStudy (campbell_st):

ok... so the width is w then the length is 5w so the perimeter is P = 2(l) + 2(w) or making the length substitution P = 2(5w) +2( w) the term at most.... to me, means less than or equal to so I would have 2w + 2(5w) < = 96

OpenStudy (anonymous):

@campbell_st A?

OpenStudy (campbell_st):

that would be my answer...

OpenStudy (anonymous):

The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 40 cm B. 19.2 cm C. 16 cm D. 8 cm